Yield to Maturity

 Suppose that a 6.250% semi-annual coupon paying bond is priced at 105.760 per 100 of par value. The tenor of the bond is 13 years.  What is the yield to maturity?
Answer #1

No of periods \(=12\) years \(^{*} 2=24\) semi annual periods

Coupon per period = (Coupon rate / No of coupon payments per year) * Face value

Coupon per period \(=(6.250 \% / 2) * \$ 100\)

Coupon per period \(=\$ 3.125\)

Bond Price \(=\sum\) Coupon \(/(1+\mathrm{YTM})\) period \(+\) Face value \(/(1+\mathrm{YTM})^{\text {period }}\)

\(\$ 105.5=\$ 3.125 /(1+\mathrm{YTM} / 2)^{1}+\$ 3.125 /\left(1+\mathrm{YTM} / 2 \mathrm{e}^{2}+\ldots+\$ 3.125 /(1+\mathrm{YTM} / 2 \mathrm{e})^{24}+\$ 100 /(1+\mathrm{YTM} /\right. 2)^{24}\)

Using Texas Instruments BA 2 Plus Calculator

\(\mathrm{SET} \mathrm{N}=24, \mathrm{FV}=100, \mathrm{PMT}=3.125, \mathrm{PV}=-105.5\)

\(\mathrm{CPT}  I / Y =2.8070\)

\(Y T M=2^{*} I / Y\)

\(\mathrm{YTM}=2 * 2.8070\)

\(\mathrm{YTM}=5.6139 \%\) or \(5.613 \%\)

answered by: Bloggeron